I need help on this problem that I have been working on forever. Every time i tried it i get a different answer each time. Here's the Problem
{2x - 3/4y = 17
{-3x - 5/8y = -11
Its an system of equations and i need to find out whether its a consistant independent, consistant dependent(infinite solution), or inconsistant(no solution) i tried solving it using the combination elimation method but im getting answers that look like this x= 116/15 and y= 113/8 its weird. so PLEASE help me. THANK YOU!!!


Answer by venugopalramana(3286)

{2x - 3/4y = 17
{-3x - 5/8y = -11
Its an system of equations and i need to find out whether its a consistant independent, consistant dependent(infinite solution), or inconsistant(no solution) i tried solving it using the combination elimation method but im getting answers that look like this x= 116/15 and y= 113/8 its weird.
DONT WORRY ABOUT ANSWER LOOKING WEIRD..THEY COME OUT AS PER THE GIVEN PROBLEM..
LET US CHECK.....................
2x - (3/4)y = 17..........................I
-3x - (5/8)y = -11 ........................II
EQN.I*3+EQN.II*2...
-(9/4)Y-(5/4)Y=51-22=29
-(14/4)Y=29
Y=-4*29/14=-58/7
SUBSTITUTING IN EQN.I..
2X+(3*58/4*7)=17
2X=17-87/14=(238-87)/14=151/14
X=151/28
SO THE ANSWERS LOOK WEIRD AS YOU SAY.BUT THEY ARE CORRECT.IT IS BECAUSE THE GIVEN PROBLEM IS OF THAT TYPE.
I DONT KNOW HOW YOU GOT YOUR ANSWER.IF YOU SEND YOUR WORKING I CAN CHECK AND TELL WHERE YOU MADE THE MISTAKE..AS I SAID BEFORE DONT WORRY ABOUT ANSWER LOOKING WEIRD..THEY COME OUT AS PER THE GIVEN PROBLEM..